# Quantum teleportation implies symmetry-protected topological order

### v2

@@ Appendix H Proof of Lemma 10
$\left[A_{1},\,Z_{f}A_{2}\right]=\left[A_{1},\,Z_{f}\right]\,A_{2}+Z_{f}\,\left[A_{1},\,A_{2}\right]=0\,.~{}~{}$ (H.137)
We note that $A_{1}$ only fails to commute with an operator on site $f$ if, in the bare protocol $\mathcal{W}^{\prime}$ (which need not be in canonical form), the single-qubit observable $A_{1}^{\prime}$ either (i) is supported only on site $f$ or (ii) is supported on some site $j\neq f$ where both $j$ and $f$ lie in the support of some physical unitary $U$ that acts after $\mathcal{M}_{1}^{\prime}$ (in the Schrödinger picture). In either case, $A_{1}^{\prime}$ lies in the light cone of site $f$ generated by time evolution alone. Hence, attaching such an operator to any logical operator $\sigma^{\nu}_{\rm L}$ cannot be more useful to teleportation than time evolution alone, which obeys $L\leq vT$ (28). However, Prop. 9 establishes that the measurement of $A_{2}$ cannot lead to $L>vT$ on its own.
Hence, both $A_{1}$ and $A_{2}$ must act trivially on $f$ (in canonical form or otherwise), so that Eq. H.137 becomes
$\left[A_{1},\,A_{2}\right]=0\,,~{}~{}$ (41)
meaning that two observables $A_{1}$ and $A_{2}$ are mutually compatible if they commute. This also implies that the channels $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ commute, so their order in Eq. H.130 is inconsequential. Moreover, all of the arguments above hold if we replace $X_{{\rm L}}$ and $Z_{{\rm L}}$ with any pair of logical operators. Note that Eq. 41 must hold in the bare protocol and in canonical form (H.130).
@@
$\mathcal{W}=\Big{[}\prod\limits_{n=1}^{k}\,\mathcal{R}_{n,x}\,\mathcal{R}_{n,x}\Big{]}\,\mathcal{M}_{2k}\,\cdots\mathcal{M}_{1}\,\mathcal{U}\,,~{}$ (H.138)
acting on logical operators, where the order of the effective recovery channels $\mathcal{R}_{n,\nu}$ (7) is unimportant, and both $\mathcal{M}$ and $\mathcal{U}$ are unchanged compared to the true protocol.
We next repeat the analysis used in the $k=1$ case, noting that the combination of $\mathcal{R}$ and $\mathcal{M}$ attaches observables to logical operators, per Lemma 8. By Prop. 9, the same observable cannot be attached to both logical basis operators for a given logical qubit $n$. Moreover, the $k=1$ case above establishes that the attached operator $\bar{A}_{s}$ (the measured observable) evolves under $\mathcal{U}$ into the initial logical operator $\sigma^{\nu}_{i_{n}}$ (and some stabilizer element). Hence, if the same operator $\bar{A}_{s}$ is attached to logical operators for the logical qubits $n$ and $n^{\prime}$, then at least one of the logical operators $n$ and $n^{\prime}$ acts nontrivially on both $i_{n}$ and $i_{n^{\prime}}$, violating Eq. 1. Hence, we assign a unique observable $A_{s}$ to each logical basis operator.
The first measured observable encountered in the Heisenberg picture is $A_{2k}$; by Lemma 8, we have that
$\left[A_{2k},\,\sigma^{\nu}_{f_{n}}\right]=0~{}~{}\forall\,n,\nu\,,~{}~{}$ (H.139)
and since the label $n$ is meaningless, suppose that $\bar{A}_{2k}$ is attached to $X_{f_{k}}$ (i.e., $k$th logical $X$ operator).
We next encounter the observable $A_{2k-1}$, finding
$\left[A_{2k-1},\,X_{f_{k}}\,A_{2k}\right]=0\,,~{}~{}$ (H.140)
and note that $A_{2k-1}$ commutes with all other logical operators. Following Eq. H.137, this implies that
$\left[A_{2k-1},\,A_{2k}\right]=0\,,~{}~{}$ (H.141)
in analogy to Eq. 41. The logical operator to which $A_{2k-1}$ is attached is unimportant, as long as each measured observable is attached to a unique logical operator.
As a result, every subsequent observable obeys one or more expressions like Eq. H.140 by Lemma 8. Since the measured observables always commute with the logical operators (which is required for $L>vT$), the observable $A_{2k-s}$ obeys Eq. H.140 for the $s$ logical operators to which prior observables have been attached, and commutes with all other logical operators. Each of the $s$ expressions of the form of Eq. H.140 reduce to Eq. H.141 for each of the measured observables already encountered. Repeating this for all $2k$ measurement channels, we find that
$\left[A_{j},\,A_{j^{\prime}}\right]=0~{}~{}\forall\,j,j^{\prime}\,,~{}~{}$ (H.142)
@@ Appendix I Proof of Lemma 11
$=\mathcal{U}^{\dagger}\,A_{\nu,j}\,\sigma^{\nu}_{f}\,\mathcal{U}=\sigma^{\nu}_{i}\,\mathsf{S}_{\nu}\,,~{}~{}$ (I.143)
where $\sigma^{\nu}_{\rm L}(T)$ is physical part of the $\nu$-type logical operator acting on the initial state, which we take to be a product state, the final equality follows from Eq. 1, and Lemma 10 guarantees that exactly one of the two observables is attached to the $\sigma^{\nu}_{\rm L}$, which we denote by
$A_{\nu,j}\equiv U_{\nu}\,\bar{A}_{\nu j}^{\prime}\,U^{\dagger}_{\nu}\,,~{}~{}$ (32)
where $\bar{A}_{j}^{\prime}$ is the involutory part (4) of the single-qubit observable $A_{j}$ acting on qubit $j$ (which depends on $\nu$), and $U_{\nu}$ comprises all physical unitaries $U$ after the measurement of $A_{\nu j}^{\prime}$ in the naïve protocol $\mathcal{W}^{\prime}$, in the Schrödinger picture. We can then write $\mathcal{U}=U_{\nu}\,\mathcal{U}_{\nu}$, so that $\mathcal{U}_{\nu}$ comprises all physical unitaries in $\mathcal{W}^{\prime}$ prior to the measurement of $A_{\nu j}^{\prime}$, where $U_{\nu}$ has depth $\tau_{\nu}$, $\mathcal{U}_{\nu}$ has depth $T_{\nu}$, and $T=\tau_{\nu}+T_{\nu}$. We then rewrite Eq. I as
$\sigma^{\nu}_{i}\,\mathsf{S}_{\nu}=\left(\mathcal{U}^{\dagger}_{\nu}\,\bar{A}_{\nu j}^{\prime}\,\mathcal{U}_{\nu}\right)\,\left(\mathcal{U}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{U}\right)\,,~{}~{}$ (I.144)
@@
Next, we establish by contradiction that the light cones of $\bar{A}_{\nu j}^{\prime}$ and $\sigma^{\nu}_{f}$ in Eq. I.144 have to meet. If the light cones do not meet, then since $\Sigma^{\nu}_{\ell}$ cannot realize $\sigma^{\nu}_{i}$, it must realize part of the stabilizer element $\mathsf{S}_{\nu}$ in Eq. I.144, while $\mathcal{U}^{\dagger}_{\nu}\,\bar{A}_{\nu j}^{\prime}\,\mathcal{U}_{\nu}$ realizes $\sigma^{\nu}_{i}$ times another part of $\mathsf{S}_{\nu}$, so that
$\mathsf{S}_{\nu}$ $=\mathsf{S}^{(1)}_{\nu}\otimes\mathsf{S}^{(2)}_{\nu}=\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}_{\nu}\,\bar{A}_{\nu j}^{\prime}\,\mathcal{U}_{\nu}\otimes\Sigma^{\nu}_{\ell}\,,~{}~{}$ (I.146)
where $\mathsf{S}^{(1)}_{\nu},\mathsf{S}^{(2)}_{\nu}\in\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$ (3) have disjoint support.
Note that any two elements $\mathsf{S}_{1},\mathsf{S}_{2}$ of $\operatorname*{stab}\left(*\right){\lvert\Psi\rangle}$ satisfy
$\lVert\left\{\mathsf{S}_{1},\,\mathsf{S}_{2}\right\}\rVert=2\,,~{}~{}$ (I.147)
which we prove for all $\mathsf{S}_{1},\mathsf{S}_{2}\in\operatorname*{stab}\left(*\right){\lvert\Psi\rangle}$. First, note that $\left\{\mathsf{S}_{1},\,\mathsf{S}_{2}\right\}\,\lvert\Psi\rangle=2\,\lvert\Psi\rangle$, which implies that $\lVert\left\{\mathsf{S}_{1},\,\mathsf{S}_{2}\right\}\rVert\geq 2$ (since there exists an eigenvalue of the anticommutator with eigenvalue at least two). Second, note that unitarity of the stabilizer elements—along with the triangle inequality—ensure that $\lVert\left\{\mathsf{S}_{1},\,\mathsf{S}_{2}\right\}\rVert\leq 2\,\lVert\mathsf{S}_{1}\,\mathsf{S}_{2}\rVert=2$.
Crucially, the initial state $\lvert\Psi_{0}\rangle$ (6) is either a product state, or is separable with respect to the support of the two stabilizer elements. However, Eq. I.146 implies that
$\left\{\mathsf{S}^{(2)}_{\mu},\,\mathsf{S}^{(2)}_{\nu}\right\}=\mathcal{U}^{\dagger}\,\left\{\sigma^{\mu}_{f},\,\sigma^{\nu}_{f}\right\}\,\mathcal{U}=2\,\delta_{\mu,\nu}\,,~{}~{}$ (I.148)
which is incompatible with Eq. I.147 for $\mu\neq\nu$, contradicting the claim that $\mathsf{S}^{(2)}_{\nu}$ are elements of $\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$ (3). Hence, the light cones of the operators $\bar{A}_{\nu j}^{\prime}$ and $\sigma^{\nu}_{f}$ must meet.
Moreover, the operators $\sigma^{\nu}_{\rm L}(T)$ (I) must realize the Pauli algebra on site $i$ only, and commute everywhere else. The most efficient strategy to ensure this is to evolve the commuting operators $A_{\nu j}^{\prime}$ (for $\nu=x,z$) to nearby sites (using a circuit with depth $\tau_{0}$, and apply a unitary “decoding” channel $\mathcal{Q}$ that maps, e.g.,
@@
As established in Ref. 36, the most efficient strategy is to evolve $\sigma^{\nu}_{j}\to\sigma^{\nu}_{i}$ and $\sigma^{\nu}_{j+v}\to\sigma^{\nu}_{\ell-v}$, with
@@
so that $\sigma^{\nu}_{\ell}\,\sigma^{\nu}_{\ell-v}$ can be converted into stabilizers of $\lvert\Psi_{0}\rangle$ (6) for both $\nu=x,z$ by the “encoding” channel $\mathcal{Q}^{\dagger}$.
@@
$j-i$ $\leq v\,\left(T-\max_{\nu}\tau_{\nu}-\tau_{0}-1\right)$ (I.152a)
$\ell-j$ $\leq v\left(T-\max_{\nu}\tau_{\nu}-\tau_{0}\right)\,,~{}~{}$ (I.152b)
where $\max_{\nu}\tau_{\nu}$ is the total depth of all physical unitary channels following the first measurement (in the Schrödinger picture). By summing the foregoing inequalities, we obtain the following bound on $L$,
$L=d(i,f)\leq 3vT-2v\left(\max_{\nu}\tau_{\nu}+\tau_{0}+1\right)\,,$ (I.153)
where the depth $\tau_{0}\geq 0$ increases if the measurements of $A_{\nu,j}^{\prime}$ are either (i) far from one another (i.e., $d(j_{1},j_{2})\gg 1$) or (ii) far from the optimal position $j^{*}\approx i+T/3$. Taking $\tau_{0}\to 0$ in Eq. I.153 gives the (potentially loose) bound
$L=d(i,f)\leq 3vT-2v\left(\max_{\nu}\tau_{\nu}+1\right)\,.~{}$ (43)
@@
As in the $M=2$ case described by Lemma 12, we need only consider the optimal scenario. This means that pairs of measurements are well separated, and all of the measured observables commute with one another and all logical operators. This applies both to the canonical-form observables $A_{j}$ (32) and the single-qubit observables $A_{j}^{\prime}$ of the naïve protocol $\mathcal{W}^{\prime}$. Evolving the logical through the dilated channels $\mathcal{R}$ and $\mathcal{M}$ leads to
$\sigma^{\nu}_{f}(t)=\sigma^{\nu}_{f}\,\prod\limits_{s=1}^{M}\,U_{s,\nu}\,\bar{A}_{s,\nu}\,U^{\dagger}_{s,\nu}\,,~{}~{}$ (J.154)
where $s$ runs over pairs of measurements (one of which is assigned to $X_{{\rm L}}$ and the other to $Z_{{\rm L}}$), and $U_{s,\nu}$ captures all unitary evolution following (in the Schrödinger picture) the measurement of the single-qubit operator $A_{s,\nu}$ in the naïve protocol $\mathcal{W}^{\prime}$ (and has depth $\tau_{s,\nu}$).
@@
as before. Again, it is optimal to measure nearby sites—e.g., $j_{s,z}=j_{s,x}+v$, as in Fig. 2 and Eq. I.149. We denote by $j_{s}$ the left site of the pair $s$. As in Lemma 12, we have
$d(j_{1},i)$ $\leq v\,\left(T-\max_{\nu}\tau_{1,\nu}-1\right)$ (I.152a)
$d(\ell,j_{m})$ $\leq v\left(T-\max_{\nu}\tau_{m,\nu}-1\right)\,,~{}~{}$ (I.152b)
where $m=M/2$ is the number of pairs of nearby measurements, and we take $\tau_{0}=0$ compared to Eq. I. The new ingredient compared to Lemma 12 is
$d(j_{s+1},j_{s})\leq v\left(T-\max_{\nu}\max(\tau_{s,\nu},\tau_{s+1,\nu})-1\right)\,,$ (J.155)
and adding these all up—and taking $\tau_{s,\nu}=0$ to reflect optimal protocols—we find that
@@
To prove Eq. 47, we derive the optimal spacing of measurements for a given logical qubit $n$, as in the proofs of Lemmas 11 and 12. This is straightforward, as Lemmas 7 and 8 jointly guarantee that the combination of measurements and outcome-dependent operations do not “mix” between logical operators. Moreover, Lemma 10—whose proof in App. H applies to all $k\geq 1$—guarantees that all of the measured observables $A_{j}$ (32)—and their involutory parts $\bar{A}_{j}$ (4)—commute with one another and with all final-state logical operators, by Lemma 8.
Evolving all logical operators through the channels $\mathcal{R}$ and $\mathcal{M}$ in the Heisenberg picture, projecting onto the default Stinespring initial state $\lvert\boldsymbol{0}\rangle_{\rm ss}$, and tracing out the Stinespring degrees of freedom leads to

### v3

@@ Appendix H Proof of Lemma 10
$\left[A^{\vphantom{\prime}}_{1},\,Z^{\vphantom{*}}_{f}A^{\vphantom{\prime}}_{2}\right]=\left[A^{\vphantom{\prime}}_{1},\,Z^{\vphantom{*}}_{f}\right]\,A^{\vphantom{\prime}}_{2}+Z^{\vphantom{*}}_{f}\,\left[A^{\vphantom{\prime}}_{1},\,A^{\vphantom{\prime}}_{2}\right]=0\,.~{}~{}$ (H.137)
We note that $A^{\vphantom{\prime}}_{1}$ only fails to commute with an operator on site $f$ if, in the bare protocol $\mathcal{W}^{\prime}$ (which need not be in canonical form), the single-qubit observable $A_{1}^{\prime}$ either (i) is supported only on site $f$ or (ii) is supported on some site $j\neq f$ where both $j$ and $f$ lie in the support of some physical unitary $U$ that acts after $\mathcal{M}_{1}^{\prime}$ (in the Schrödinger picture). In either case, $A_{1}^{\prime}$ lies in the light cone of site $f$ generated by time evolution alone. Hence, attaching such an operator to any logical operator $\sigma^{\nu}_{\rm L}$ cannot be more useful to teleportation than time evolution alone, which obeys $L\leq vT$ (28). However, Prop. 9 establishes that the measurement of $A^{\,}_{2}$ cannot lead to $L>vT$ on its own.
Hence, both $A^{\,}_{1}$ and $A^{\,}_{2}$ must act trivially on $f$ (in canonical form or otherwise), so that Eq. H.137 becomes
$\left[A^{\vphantom{\prime}}_{1},\,A^{\vphantom{\prime}}_{2}\right]=0\,,~{}~{}$ (41)
meaning that two observables $A^{\,}_{1}$ and $A^{\,}_{2}$ are mutually compatible if they commute. This also implies that the channels $\mathcal{M}^{\,}_{1}$ and $\mathcal{M}^{\,}_{2}$ commute, so their order in Eq. H.130 is inconsequential. Moreover, all of the arguments above hold if we replace $X^{\vphantom{*}}_{{\rm L}}$ and $Z^{\vphantom{*}}_{{\rm L}}$ with any pair of logical operators. Note that Eq. 41 must hold in the bare protocol and in canonical form (H.130).
@@
$\mathcal{W}=\Big{[}\prod\limits_{n=1}^{k}\,\mathcal{R}^{\vphantom{\dagger}}_{n,x}\,\mathcal{R}^{\vphantom{\dagger}}_{n,x}\Big{]}\,\mathcal{M}^{\vphantom{\dagger}}_{2k}\,\cdots\mathcal{M}^{\vphantom{\dagger}}_{1}\,\mathcal{U}\,,~{}$ (H.138)
acting on logical operators, where the order of the effective recovery channels $\mathcal{R}^{\,}_{n,\nu}$ (38) is unimportant, and both $\mathcal{M}$ and $\mathcal{U}$ are unchanged compared to the true protocol.
We next repeat the analysis used in the $k=1$ case, noting that the combination of $\mathcal{R}$ and $\mathcal{M}$ attaches observables to logical operators, per Lemma 8. By Prop. 9, the same observable cannot be attached to both logical basis operators for a given logical qubit $n$. Moreover, the $k=1$ case above establishes that the attached operator $\bar{A}^{\,}_{s}$ (the measured observable) evolves under $\mathcal{U}$ into the initial logical operator $\sigma^{\nu}_{i_{n}}$ (and some stabilizer element). Hence, if the same operator $\bar{A}^{\,}_{s}$ is attached to logical operators for the logical qubits $n$ and $n^{\prime}$, then at least one of the logical operators $n$ and $n^{\prime}$ acts nontrivially on both $i_{n}$ and $i_{n^{\prime}}$, violating Eq. 8. Hence, we assign a unique observable $A^{\,}_{s}$ to each logical basis operator.
The first measured observable encountered in the Heisenberg picture is $A^{\,}_{2k}$; by Lemma 8, we have that
$\left[A^{\vphantom{\prime}}_{2k},\,\sigma^{\nu}_{f_{n}}\right]=0~{}~{}\forall\,n,\nu\,,~{}~{}$ (H.139)
and since the label $n$ is meaningless, suppose that $\bar{A}^{\,}_{2k}$ is attached to $X^{\vphantom{*}}_{f_{k}}$ (i.e., $k$th logical $X^{\vphantom{*}}$ operator).
We next encounter the observable $A^{\,}_{2k-1}$, finding
$\left[A^{\vphantom{\prime}}_{2k-1},\,X^{\vphantom{*}}_{f_{k}}\,A^{\vphantom{\prime}}_{2k}\right]=0\,,~{}~{}$ (H.140)
and note that $A^{\,}_{2k-1}$ commutes with all other logical operators. Following Eq. H.137, this implies that
$\left[A^{\vphantom{\prime}}_{2k-1},\,A^{\vphantom{\prime}}_{2k}\right]=0\,,~{}~{}$ (H.141)
in analogy to Eq. 41. The logical operator to which $A^{\,}_{2k-1}$ is attached is unimportant, as long as each measured observable is attached to a unique logical operator.
As a result, every subsequent observable obeys one or more expressions like Eq. H.140 by Lemma 8. Since the measured observables always commute with the logical operators (which is required for $L>vT$), the observable $A^{\,}_{2k-s}$ obeys Eq. H.140 for the $s$ logical operators to which prior observables have been attached, and commutes with all other logical operators. Each of the $s$ expressions of the form of Eq. H.140 reduce to Eq. H.141 for each of the measured observables already encountered. Repeating this for all $2k$ measurement channels, we find that
$\left[A^{\vphantom{\prime}}_{j},\,A^{\vphantom{\prime}}_{j^{\prime}}\right]=0~{}~{}\forall\,j,j^{\prime}\,,~{}~{}$ (H.142)
@@ Appendix I Proof of Lemma 11
$=\mathcal{U}^{\dagger}\,A^{\vphantom{\prime}}_{\nu,j}\,\sigma^{\nu}_{f}\,\mathcal{U}=\sigma^{\nu}_{i}\,\mathsf{S}^{\vphantom{\prime}}_{\nu}\,,~{}~{}$ (I.143)
where $\sigma^{\nu}_{\rm L}(T)$ is the physical part of the $\nu$-type logical operator acting on the initial state and the final equality follows from Eq. 8. Lemma 10 guarantees that exactly one of the two measured observables is attached to $\sigma^{\nu}_{\rm L}$; the attached operator is given by
$A^{\vphantom{\prime}}_{\nu,j}\equiv U^{\vphantom{\dagger}}_{\nu}\,\bar{A}_{\nu j}^{\prime}\,U^{\dagger}_{\nu}\,,~{}~{}$ (32)
where $\bar{A}_{j}^{\prime}$ is the involutory part (33) of the single-qubit observable $A^{\,}_{j}$ acting on qubit $j$ (which depends on $\nu$), and $U^{\,}_{\nu}$ comprises all physical unitaries $U$ after the measurement of $A_{\nu j}^{\prime}$ in the naïve protocol $\mathcal{W}^{\prime}$, in the Schrödinger picture. We can then write $\mathcal{U}=U^{\,}_{\nu}\,\mathcal{U}^{\,}_{\nu}$, so that $\mathcal{U}^{\,}_{\nu}$ comprises all physical unitaries in $\mathcal{W}^{\prime}$ prior to the measurement of $A_{\nu j}^{\prime}$, where $U^{\,}_{\nu}$ has depth $\tau^{\,}_{\nu}$, $\mathcal{U}^{\,}_{\nu}$ has depth $T^{\,}_{\nu}$, and $T=\tau^{\,}_{\nu}+T^{\,}_{\nu}$. We then rewrite Eq. I as
$\sigma^{\nu}_{i}\,\mathsf{S}^{\vphantom{\prime}}_{\nu}=\left(\mathcal{U}^{\dagger}_{\nu}\,\bar{A}_{\nu j}^{\prime}\,\mathcal{U}^{\vphantom{\dagger}}_{\nu}\right)\,\left(\mathcal{U}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{U}\right)\,,~{}~{}$ (I.144)
@@
Next, we establish by contradiction that the light cones of $\bar{A}_{\nu j}^{\prime}$ and $\sigma^{\nu}_{f}$ in Eq. I.144 have to meet. If the light cones do not meet, then since $\Sigma^{\nu}_{\ell}$ cannot realize $\sigma^{\nu}_{i}$, it must realize part of the stabilizer element $\mathsf{S}^{\,}_{\nu}$ in Eq. I.144, while $\mathcal{U}^{\dagger}_{\nu}\,\bar{A}_{\nu j}^{\prime}\,\mathcal{U}^{\vphantom{\dagger}}_{\nu}$ realizes $\sigma^{\nu}_{i}$ times another part of $\mathsf{S}^{\,}_{\nu}$, so that
$\mathsf{S}^{\vphantom{\prime}}_{\nu}$ $=\mathsf{S}^{(1)}_{\nu}\otimes\mathsf{S}^{(2)}_{\nu}=\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}_{\nu}\,\bar{A}_{\nu j}^{\prime}\,\mathcal{U}^{\vphantom{\dagger}}_{\nu}\otimes\Sigma^{\nu}_{\ell}\,,~{}~{}$ (I.146)
where $\mathsf{S}^{(1)}_{\nu},\mathsf{S}^{(2)}_{\nu}\in\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)$ (3) have disjoint support.
Note that any two elements $\mathsf{S}^{\,}_{1},\mathsf{S}^{\,}_{2}$ of $\operatorname*{stab}\left(*\right){\lvert\Psi\rangle}$ satisfy
$\lVert\left\{\mathsf{S}^{\vphantom{*}}_{1},\,\mathsf{S}^{\vphantom{*}}_{2}\right\}\rVert=2\,,~{}~{}$ (I.147)
which we prove for all $\mathsf{S}^{\,}_{1},\mathsf{S}^{\,}_{2}\in\operatorname*{stab}\left(*\right){\lvert\Psi\rangle}$. First, note that $\left\{\mathsf{S}^{\,}_{1},\,\mathsf{S}^{\,}_{2}\right\}\,\lvert\Psi\rangle=2\,\lvert\Psi\rangle$, which implies that $\lVert\left\{\mathsf{S}^{\,}_{1},\,\mathsf{S}^{\,}_{2}\right\}\rVert\geq 2$ (since there exists an eigenvalue of the anticommutator with eigenvalue at least two). Second, note that unitarity of the stabilizer elements—along with the triangle inequality—ensure that $\lVert\left\{\mathsf{S}^{\,}_{1},\,\mathsf{S}^{\,}_{2}\right\}\rVert\leq 2\,\lVert\mathsf{S}^{\,}_{1}\,\mathsf{S}^{\,}_{2}\rVert=2$.
Crucially, the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6) is either a product state, or is separable with respect to the support of the two stabilizer elements. However, Eq. I.146 implies that
$\left\{\mathsf{S}^{(2)}_{\mu},\,\mathsf{S}^{(2)}_{\nu}\right\}=\mathcal{U}^{\dagger}\,\left\{\sigma^{\mu}_{f},\,\sigma^{\nu}_{f}\right\}\,\mathcal{U}=2\,\delta^{\,}_{\mu,\nu}\,,~{}~{}$ (I.148)
which is incompatible with Eq. I.147 for $\mu\neq\nu$, contradicting the claim that $\mathsf{S}^{(2)}_{\mu,\nu}$ are elements of $\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)$ (3). Hence, the light cones of the operators $\bar{A}_{\nu j}^{\prime}$ and $\sigma^{\nu}_{f}$ must meet.
Moreover, the operators $\sigma^{\nu}_{\rm L}(T)$ (I) must realize the Pauli algebra on site $i$ only, and commute everywhere else. The most efficient strategy to ensure this is to evolve the commuting operators $A_{\nu j}^{\prime}$ (for $\nu=x,z$) to nearby sites (using a circuit with depth $\tau^{\,}_{0}$, and apply a unitary “decoding” channel $\mathcal{Q}$ that maps, e.g.,
@@
As established in Ref. 42, the most efficient strategy is to evolve $\sigma^{\nu}_{j}\to\sigma^{\nu}_{i}$ and $\sigma^{\nu}_{j+v}\to\sigma^{\nu}_{\ell-v}$, with
@@
so that $\sigma^{\nu}_{\ell}\,\sigma^{\nu}_{\ell-v}$ can be converted into stabilizers of $\lvert\Psi^{\,}_{0}\rangle$ (6) for both $\nu=x,z$ by the “encoding” channel $\mathcal{Q}^{\dagger}$.
@@
$j-i$ $\leq v\,\left(T-\max_{\nu}\tau^{\vphantom{\prime}}_{\nu}-\tau^{\vphantom{\prime}}_{0}-1\right)$ (I.152a)
$\ell-j$ $\leq v\left(T-\max_{\nu}\tau^{\vphantom{\prime}}_{\nu}-\tau^{\vphantom{\prime}}_{0}\right)\,,~{}~{}$ (I.152b)
where $\max_{\nu}\tau^{\,}_{\nu}$ is the total depth of all physical unitary channels following the first measurement (in the Schrödinger picture). By summing the foregoing inequalities, we obtain the following bound on $L$,
$L=d(i,f)\leq 3vT-2v\left(\max_{\nu}\tau^{\,}_{\nu}+\tau^{\vphantom{\prime}}_{0}+1\right)\,,$ (I.153)
where the depth $\tau^{\,}_{0}\geq 0$ increases if the measurements of $A_{\nu,j}^{\prime}$ are either (i) far from one another (i.e., $d(j^{\,}_{1},j^{\,}_{2})\gg 1$) or (ii) far from the optimal position $j^{*}\approx i+T/3$. Taking $\tau^{\,}_{0}\to 0$ in Eq. I.153 gives the (potentially loose) bound
$L=d(i,f)\leq 3vT-2v\left(\max_{\nu}\tau^{\vphantom{\prime}}_{\nu}+1\right)\,.~{}$ (43)
@@
As in the $M=2$ case described by Lemma 12, we need only consider the optimal scenario. This means that pairs of measurements are well separated, and all of the measured observables commute with one another and all logical operators. This applies both to the canonical-form observables $A^{\,}_{j}$ (32) and the single-qubit observables $A_{j}^{\prime}$ of the naïve protocol $\mathcal{W}^{\prime}$. Evolving the logical through the dilated channels $\mathcal{R}$ and $\mathcal{M}$ leads to
$\sigma^{\nu}_{f}(t)=\sigma^{\nu}_{f}\,\prod\limits_{s=1}^{M}\,U^{\vphantom{\dagger}}_{s,\nu}\,\bar{A}^{\vphantom{\dagger}}_{s,\nu}\,U^{\dagger}_{s,\nu}\,,~{}~{}$ (J.154)
where $s$ runs over pairs of measurements (one of which is assigned to $X^{\vphantom{*}}_{{\rm L}}$ and the other to $Z^{\vphantom{*}}_{{\rm L}}$), and $U^{\,}_{s,\nu}$ captures all unitary evolution following (in the Schrödinger picture) the measurement of the single-qubit operator $A_{s,\nu}$ in the naïve protocol $\mathcal{W}^{\prime}$ (and has depth $\tau^{\,}_{s,\nu}$).
@@
as before. Again, it is optimal to measure nearby sites—e.g., $j^{\,}_{s,z}=j^{\,}_{s,x}+v$, as in Fig. 2 and Eq. I.149. We denote by $j^{\,}_{s}$ the left site of the pair $s$. As in Lemma 12, we have
$d(j^{\vphantom{\prime}}_{1},i)$ $\leq v\,\left(T-\max_{\nu}\tau^{\vphantom{\prime}}_{1,\nu}-1\right)$ (I.152a)
$d(\ell,j^{\vphantom{\prime}}_{m})$ $\leq v\left(T-\max_{\nu}\tau^{\vphantom{\prime}}_{m,\nu}-1\right)\,,~{}~{}$ (I.152b)
where $m=M/2$ is the number of pairs of nearby measurements, and we take $\tau^{\vphantom{\prime}}_{0}=0$ compared to Eq. I.152. The new ingredient compared to Lemma 12 is
$d(j^{\vphantom{\prime}}_{s+1},j^{\vphantom{\prime}}_{s})\leq v\left(T-\max_{\nu}\max(\tau^{\vphantom{\prime}}_{s,\nu},\tau^{\vphantom{\prime}}_{s+1,\nu})-1\right)\,,$ (J.155)
and adding these all up—and taking $\tau^{\,}_{s,\nu}=0$ to reflect optimal protocols—we find that
@@
To prove Eq. 47, we derive the optimal spacing of measurements for a given logical qubit $n$, as in the proofs of Lemmas 11 and 12. This is straightforward, as Lemmas 7 and 8 jointly guarantee that the combination of measurements and outcome-dependent operations do not “mix” between logical operators. Moreover, Lemma 10—whose proof in App. H applies to all $k\geq 1$—guarantees that all of the measured observables $A^{\,}_{j}$ (32)—and their involutory parts $\bar{A}^{\,}_{j}$ (33)—commute with one another and with all final-state logical operators, by Lemma 8.
Evolving all logical operators through the channels $\mathcal{R}$ and $\mathcal{M}$ in the Heisenberg picture, projecting onto the default Stinespring initial state $\lvert\boldsymbol{0}\rangle^{\,}_{\rm ss}$, and tracing out the Stinespring degrees of freedom leads to